# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Appendix H Remaining Experimental Details
Appendix I Consensus-Based Retraining Does Better than Confidence-Based Retraining
Here we compare full and consensus-based RT against another strategy for retraining which we call confidence-based retraining (RT). Specifically, we propose to retrain with the predicted labels of the samples with the top $50\%$ margin (i.e., highest predicted probability - second highest predicted probability); margin is a measure of the model’s confidence. This idea is similar to self-training’s method of sample selection in the semi-supervised setting (Amini et al.,, 2022). In Tables 7 and 8, we show results for CIFAR-10 and CIFAR-100 (in the same setting as Section 5 and Appendix H) with the smallest value of $\epsilon$ from Tables 1 and 2, respectively. Notice that consensus-based RT is clearly better than confidence-based RT.
@@ Appendix I Consensus-Based Retraining Does Better than Confidence-Based Retraining
Even though our empirical focus in this paper has been label DP training, retraining (RT) can be employed for general problems with label noise. Here we evaluate RT in a setting with “real” label noise due to human annotation. Specifically, we focus on training a ResNet-18 model (without label DP to be clear) on the CIFAR-100N dataset introduced by Wei et al., (2021) and available on the TensorFlow website. CIFAR-100N is just CIFAR-100 labeled by humans; thus, it has real human annotation errors. The experimental setup and details are the same as CIFAR-100 (as stated in Section 5 and Appendix H); the only difference is that here we use initial learning rate $=0.01$.

### v2

@@ Appendix G Remaining Experimental Details
Appendix H Consensus-Based Retraining Does Better than Confidence-Based Retraining
Here we compare full and consensus-based RT against another strategy for retraining which we call confidence-based retraining (RT). Specifically, we propose to retrain with the predicted labels of the samples with the top $50\%$ margin (i.e., highest predicted probability - second highest predicted probability); margin is a measure of the model’s confidence. This idea is similar to self-training’s method of sample selection in the semi-supervised setting (Amini et al.,, 2022). In Tables 8 and 9, we show results for CIFAR-10 and CIFAR-100 (in the same setting as Section 5 and Appendix G) with the smallest value of $\epsilon$ from Tables 1 and 2, respectively. Notice that consensus-based RT is clearly better than confidence-based RT.
@@ Appendix H Consensus-Based Retraining Does Better than Confidence-Based Retraining
Appendix I Retraining is Beneficial Even Without a Validation Set
In all our previous experiments, we assumed access to a small clean validation set. Here we consider training ResNet-34 on CIFAR-100 when we do not have access to a validation set and keep training for a fixed number of epochs (100). The point of these experiments is to show that retraining can offer gains even without a validation set, in which case we can expect severe overfitting. Please see the the results and discussion in Table 10.
Table 10: No validation set (CIFAR-100 w/ ResNet-34). Test set accuracies (mean $\pm$ standard deviation). Just like all our previous results in the presence of a small validation set, consensus-based RT is the clear winner here. However, due to a higher degree of overfitting here, the performance and amount of improvement obtained with retraining is worse than the corresponding experiment using a validation set (Table 4). This is not surprising.
$\epsilon$ Baseline Full RT Consensus-based RT
3 $14.53\pm 0.48$ $15.20\pm 0.71$ $\textbf{17.30}\pm 0.43$
4 $26.03\pm 1.75$ $28.33\pm 1.76$ $\textbf{30.47}\pm 1.31$
5 $42.50\pm 1.14$ $44.43\pm 1.27$ $\textbf{46.27}\pm 1.72$
@@ Appendix I Retraining is Beneficial Even Without a Validation Set
Even though our empirical focus in this paper has been label DP training, retraining (RT) can be employed for general problems with label noise. Here we evaluate RT in a setting with “real” label noise due to human annotation. Specifically, we focus on training a ResNet-18 model (without label DP to be clear) on the CIFAR-100N dataset introduced by Wei et al., (2021) and available on the TensorFlow website. CIFAR-100N is just CIFAR-100 labeled by humans; thus, it has real human annotation errors. The experimental setup and details are the same as CIFAR-100 (as stated in Section 5 and Appendix G); the only difference is that here we use initial learning rate $=0.01$.